Home‎ > ‎Crew 136‎ > ‎

226-CommandersReport

Commander Report
02/26/2014
Corentin Liber

In the morning, the data from the balloon launch was collected and then analyzed later in the
afternoon. We are currently writing a first science report about these results.

The same went for the geology project.  A first study was concluded in the north area yesterday and
a report is now available about the conclusions of the sampling analysis.  A second area was checked
today while on ATV and some samples were again collected.  A third zone should be visited tomorrow
to conclude this project.

We also worked on the data from our streambed project and obtained some results.  We'll analyze the
rest of them by tomorrow and write a complete report about our conclusions and model.

The astronomy project is depending on the weather and won't be possible this evening if the
conditions don't improve in the following hours.  The spectrum recorded are currently being analyzed.

--------------------------

CommanderCheckIn Report
02/26/2014
Thibaud Herriau

Crew Physical Health:
All are well.

Eva Depart Time:
2pm
Eva Return Time:
4:30pm
Narrative Of Field Mission Results:
Exploration new area with ATV.
Due to the scary clouds and the bad roads, the EVA subcrew had to turn
around and return to the hab before reaching the target area.
No samples collected.

Hab Maint:
Refill of the ATV fuel reservoirs.

We are out of garbage bags & sponges.
Crew137 should be notified so they can bring new ones.
Rep Sched:
None
Tomorrow Plan:
Long range EVA with ATV.
Analysis of the spectra of the moon, sun and other sky objects.
Data mining of the data of the streambeds.
Referencing all samples on ArcGIS
Inventory:
None
Support Req:
None
Miscelleneous:
Officer Herriau still has his name misspelled on
http://mdrs.marssociety.org/home/crew-136
Bottom of the page, as title of his the biography.
>> The correct spelling is "Herriau"